# Gravitational Wave Signal Extraction Against Non-Stationary Instrumental Noises with Deep Neural Network

### v2

@@ I INTRODUCTION
Deep learning methods have already achieved considerable success in data processing of GW detections [46, 47, 32, 48, 49, 50, 51, 52, 53, 54, 55, 31]. Compared to traditional algorithms, such as matched filtering [29, 30, 56], which generally require a vast template bank and thus consume substantial computational resources, deep learning methods are well known for their fast processing speed and excellent generalization ability, making it possible to achieve rapid and anomaly-model-independent signal extraction. For GW detection in space, Conv-TasNet [31] has demonstrated its performance in various scenarios, claiming an extraction speed improvement of up to times compared to traditional methods, such as the Markov-chain Monte Carlo and genetic algorithms [57, 58, 59, 60, 61], however, robustness tests under anomalous conditions are overlooked. In this work, focusing on the three types of possible non-stationarities for LISA or LISA-like missions, including gaps, transients (glitches), and time varying noise auto-correlations, we developed a deep learning model for accurate extractions of coalescing MBHB signals confronted with such anomalous scenarios. This also provide new explorations into the robustness studies of deep learning models for data processing in space-borne GW missions. Our model exhibits the same performance as the Conv-TasNet model [31] does for the ideal and anomaly free scenario, while shows remarkable adaptability in signal extractions against all three types of non-stationarities and even their mixtures.
@@ III Data preparations
For LISA and LISA-like missions, laser frequency instability noise is dominant in the science measurements, and a pre-processing technology called Time Delay Interferometry (TDI) [79, 80, 81] is employed to efficiently suppress such noise. Following this, our data is generated through the first generation and noise orthogonal TDI channels and , see [79] for details. The transformed signals from incident GWs through the and channels reads
where is the transfer function combining antenna responses and TDI transformations, and represents the polarization components of the coalescing MBHB GW signal. Specific details of this transfer function can be found in [81, 79]. The PyCBC package [82] and the waveform model SEOBNRv4 [83] are employed to generate the coalescing MBHB GW templates . We show in Fig. 2 how the template waveform will change when applying the TDI combinations. The noise floor of the output data from certain TDI channel is in principle determined by the corresponding TDI combinations of GRS residual acceleration noises and optical readout noises . Therefore, the total noise Power Spectral Density (PSD) of the , channels are the TDI combinations of the PSDs of the two noise components and , see Ref [81, 79, 84, 85] for their specific combination forms, and according to the designs of LISA [19] and Taiji [86] we assume the nominal values
The transformed waveform data from a certain channel will be superimposed with the noises that go through the same TDI channel, which will serve as the simulated inputs into our model.
We established the parameter space for coalescing MBHBs as in Table 1, and employed a uniform grid to generate the diverse waveforms. Parameters such as times and phases at coalescence, sky locations, polarization angles, as well as inclinations are drawn from uniform distributions.
@@
We inject the signal with specific optimal Signal-to-Noise Ratios (SNR)
@@
here, Hz and Hz. The tilde symbol represents the Fourier transform and the complex conjugation. We set the SNR level for the training dataset as 50, and generate 10,000 samples. The distances of the sources are adjusted according to the specific optimal SNR. All samples are subjected to whitening processing, and a Tukey window with is adopted. The overlap between the signal waveform extracted by the model from the TDI output data and the whitened template waveform through channel or can be evaluated in terms of the overlap function
@@ V.1 Data gaps
We delved further into the threshold performance of our model, as detailed in Fig 6. The model exhibits minimal accuracy loss in scenarios with scheduled gaps before mergers. When confronted with the random unscheduled gaps, the model’s results gradually worsen as the frequency of occurrences increases. As expected, signals with lower SNRs are more susceptible to data gaps. In Tab 2, we present more detailed results under different conditions. According to the predetermined criterion, our model can maintain robustness even when the frequency of unscheduled gaps reaches once per 5-hours.
@@ V.2 Time-varying noise auto-correlations
While, given the predefined criterion, one finds that noises originating from the optical metrology system has a more significant impact on the final outcomes. Further investigation shows that this is because that, increasing the optical readout noise (concentrated in the relatively high frequency band) will significantly reduce the SNR of coalescing MBHB signals due to the specific forms of the PSDs of the total noise and expected signal, see Fig 9 for illustrations. This, but not the variations in noise PSDs, will make it more difficult for the model to extract signals and produce the results shown Fig 8. and Tab 3.
@@ V.3 Glitches
where is the time of glitch injection, is the impulse transferred by the glitch (i.e., the gain of test mass velocity caused by the glitch), are time scales, and denotes the Heaviside step function. Glitches are simulated and injected to one of the test mass interferometer measurement. The detailed transfer functions of glitches to the TDI outputs can be found in Ref. [39]. We considered two types of glitches according to Ref. [40], that the short-duration glitches lasting approximately 70 seconds and long-duration glitches for about 3 hours. The initial amplitudes of both types of glitches were set to , see Tab 4 for detailed parameters, and we tested the robustness of our model by increasing the amplitudes by 0.5 times for each new experiment until the performance fell below the threshold we set.
@@
Our model proficiently extracts the coalescing MBHB signals given such instrumental transients. Fig 10 shows an example of the model’s extraction result.
@@ V.4 Mixing three types of non-stationarities
Gap every 10 hours | 1x | 2x | Long | 81.8% | 0.905 | 0.216 |
@@ VI Concluding remarks
In this work, we developed a deep learning model for accurate GW extractions against possible data anomalies or non-stationarities for space-borne GW antennas, such as LISA and LISA-like missions, which also provide robustness studies of deep learning models in data processing of space-borne GW missions. Our research focuses on the three types of non-stationarities, including gaps, glitches, and time varying noise auto-correlations. Compared with the current state-of-the-art models, our model exhibits the same performance for the ideal and anomaly free cases in signal extractions. While, confronted with the three types of non-stationarities and even their mixtures in some extreme cases, our model still shows considerable adaptability. It is worth noting that when all anomalies occur at the same time, the model’s tolerance for these anomalies would degrade compared with each stand-alone situation, and among these the unscheduled gap turns out to be a more important factors.
@@ Acknowledgements.
- Cornish and Shuman [2020] N. J. Cornish and K. Shuman, Black hole hunting with lisa, Physical Review D 101, 124008 (2020).
- Crowder et al. [2006] J. Crowder, N. J. Cornish, and J. L. Reddinger, Lisa data analysis using genetic algorithms, Physical Review D 73, 063011 (2006).
- Cornish and Porter [2007a] N. J. Cornish and E. K. Porter, Catching supermassive black hole binaries without a net, Physical Review D 75, 021301 (2007a).
- Cornish and Porter [2007b] N. J. Cornish and E. K. Porter, The search for massive black hole binaries with lisa, Classical and Quantum Gravity 24, 5729 (2007b).
- Feroz et al. [2009] F. Feroz, M. Hobson, and M. Bridges, Multinest: an efficient and robust bayesian inference tool for cosmology and particle physics, Monthly Notices of the Royal Astronomical Society 398, 1601 (2009).

### v3

@@ I INTRODUCTION
Deep learning methods have already achieved considerable success in data processing of GW detections [46, 47, 32, 48, 49, 50, 51, 52, 53, 54, 55, 31]. Compared to traditional algorithms, such as matched filtering [29, 30, 56], which generally require a vast template bank and thus consume substantial computational resources, deep learning methods are well known for their fast processing speed and excellent generalization ability, making it possible to achieve rapid and anomaly-model-independent signal extraction. For GW detection in space, Conv-TasNet [31] has demonstrated its performance in various scenarios, completing signal extraction in seconds, however, robustness tests under anomalous conditions are overlooked. In this work, focusing on the three types of possible non-stationarities for LISA or LISA-like missions, including gaps, transients (glitches), and time varying noise auto-correlations, we developed a deep learning model for accurate extractions of coalescing MBHB signals confronted with such anomalous scenarios. This also provide new explorations into the robustness studies of deep learning models for data processing in space-borne GW missions. Our model exhibits the same performance as the Conv-TasNet model [31] does for the ideal and anomaly free scenario, while shows remarkable adaptability in signal extractions against all three types of non-stationarities and even their mixtures.
@@ III Data preparations
For LISA and LISA-like missions, laser frequency instability noise is dominant in the science measurements, and a pre-processing technology called Time Delay Interferometry (TDI) [74, 75, 76] is employed to efficiently suppress such noise. Following this, our data is generated through the first generation noise orthogonal TDI channels and , see [74] for a detailed introduction. The total TDI output data stream of channel (dubbed ) is the combination of signal and noise :
where the detector response in terms of channel to some incident GW with polarizations reads
where is the total transfer function of signal including antenna responses and TDI combinations, and represents the polarization components of the coalescing MBHB GW signal. Specific details of this transfer function can be found in [76, 74]. In this paper, the TDI transfer functions of signals are derived according to Taiji’s mission concept and orbit configurations [77], which has a nominal arm length of m. The PyCBC package [78] and the waveform model SEOBNRv4 [79] are employed to generate the coalescing MBHB GW templates . On the other hand, the noise floor of the output data from a certain TDI channel is in principle determined by the corresponding TDI combinations of GRS residual acceleration noises and optical metrology system noises . The resulting total instrumental noise of TDI channel can be written as
with and being the transfer functions of GRS residual acceleration noise and optical metrology system noise, respectively [74, 76], which are common to LISA and Taiji missions. The time series of basic instrumental noises and are generated according to the Power Spectral Densities (PSDs) of these two noise components, denoted as and , and according to the designs of Taiji [77] we assume the nominal values
@@
Specifically, to generate the noise data in time domain according to the PSDs, we convert Gaussian white noises to the frequency domain through Fourier transform, then multiply them by the amplitude spectral density value at each frequency to adjust the spectra of the noises, and finally obtain time series of noises through inverse Fourier transform. Theoretically, the total noise PSDs of the , channels are the TDI combinations of and calculated according to Eq. (3), see Refs. [76, 74, 80, 81] for their specific forms.
We established the parameter space for coalescing MBHBs as in Table 1, and employed a uniform grid to generate the diverse waveforms. Parameters such as times and phases at coalescence, sky locations, polarization angles, as well as inclinations are drawn from uniform distributions. We inject the signal with specific optimal Signal-to-Noise Ratios (SNR)
@@
here, Hz and Hz. The tilde symbol represents the Fourier transform and the complex conjugation. This inner product can also be used to calculate the overlap between the signal waveform extracted by the model and the whitened template waveform through channel or can be evaluated in terms of the overlap function
@@
We set the SNR level for the training dataset as 50, and generate 10,000 samples. The distances of the sources are adjusted according to the specific optimal SNR. All samples are subjected to whitening processing, and a Tukey window with is adopted.
@@
Our training and test datasets were sampled from redshifts ranging from 1 to 20. The resulting distribution is shown in Fig 3.
@@ V.1 Data gaps
We delved further into the threshold performance of our model, as detailed in Fig 6. The model exhibits minimal accuracy loss in scenarios with scheduled gaps before mergers. When confronted with the random unscheduled gaps, the model’s results gradually worsen as the frequency of occurrences increases. As expected, signals with lower SNRs are more susceptible to data gaps.
@@
In Tab 2, we present more detailed results under different conditions. According to the predetermined criterion, our model can maintain robustness even when the frequency of unscheduled gaps reaches once per 5-hours.
@@ V.2 Time-varying noise auto-correlations
While, given the predefined criterion, one finds that noises originating from the optical metrology system has a more significant impact on the final outcomes. Further investigation shows that this is because that, increasing the optical metrology system noise (concentrated in the relatively high frequency band) will significantly reduce the SNR of coalescing MBHB signals due to the specific forms of the PSDs of the total noise and expected signal, see Fig 9 for illustrations. This, but not the variations in noise PSDs, will make it more difficult for the model to extract signals and produce the results shown Fig 7. and Tab 3.
@@ V.3 Glitches
where is the time of glitch injection, is the impulse transferred by the glitch (i.e., the gain of test mass velocity caused by the glitch), are time scales, and denotes the Heaviside step function. Glitches are simulated and injected to the GRS residual acceleration noise of TM12 (i.e. the test-mass onboard spacecraft 1 and facing spacecraft 2). They are then passed through (in Eq. (3)), together with the “normal” acceleration noises, to get the total TDI combinations with glitches. We considered two types of glitches according to Ref. [40], that the short-duration glitches lasting approximately 70 seconds and long-duration glitches for about 3 hours. The initial amplitudes of both types of glitches were set to , see Tab 4 for detailed parameters, and we tested the robustness of our model by increasing the amplitudes by 0.5 times for each new experiment until the performance fell below the threshold we set.
@@
All glitches were injected with an initial time , and the time difference between the injection points of the glitches and the mergers rang from 22 to 38 hours. Our model proficiently extracts the coalescing MBHB signals given such instrumental transients. Fig 10 shows an example of the model’s extraction result.
@@ V.4 Mixing three types of non-stationarities
| Gap every 10 hours | 1x | 2x | Long | 81.8% | 0.905 | 0.216 | ||
@@ VI Concluding remarks
In this work, we developed a deep learning model for accurate GW extractions against possible data anomalies or non-stationarities for space-borne GW antennas, such as LISA and LISA-like missions, which also provide robustness studies of deep learning models in data processing of space-borne GW missions. Our research focuses on the three types of non-stationarities, including gaps, glitches, and time varying noise auto-correlations. Compared with the current state-of-the-art models, our model exhibits the same performance for the ideal and anomaly free cases in signal extractions. While, confronted with the three types of non-stationarities and even their mixtures in some extreme cases, our model still shows considerable adaptability. It is worth noting that when all anomalies occur at the same time, the model’s tolerance for these anomalies would degrade compared with each stand-alone situation, and among these the unscheduled gap turns out to be a more important factor.